Consider the demand for a good. At price Rs 4, the demand for the good is 25 units.

Suppose price of the good increases to Rs 5, and as a result, the demand for the good

falls to 20 units. Calculate the price elasticity?


1
Expert's answer
2021-09-15T17:05:31-0400

Price elasticity of demand "(E_d)" "=(-) \frac {P} {Q} ×\frac {\Delta Q} {\Delta P}"

Given, P=Rs.4;P1=Rs.5

"\Delta P=P_1-P=Rs.5-Rs.4=Rs.1"

Q=25 units ;Q1 =20units

"\Delta Q=Q_1-Q=(20-25)units =(-) 5units"

"E_d=(-) \frac {4}{25}×\frac {-5}{1}"

"=0.8"